A 60 W, 120 V lightbulb has a resistance of 12.7 ? when cold (20.0
Answer #1

Power Consumption when it is turned on

P=V2/R =1202/12.7 =1133.86 Watts

Power Consumption when it is hot

P=1202/162 =88.89 Watts